# What is a Social Media Bot? A Global Comparison of Bot and Human Characteristics

### v1

@@ Methods
We collected a dataset from Twitter/X involving global events which provides a richness in a general understanding of the bot and human differentiation. The list of data collection parameters are detailed in Appendix Table 8.
@@
To classify identities, we compare the user description and bio information against a survey of occupations of United States users performed in 2015 [105]. If the occupation is present in the user information, the user is tagged with the identity. A user can have more than one identity. We compare the top identities used by bots and human users across all events. These identities are also divided up into seven categories: religion, race/nationality, political, job, gender, family and others. We then classify each user into these categories of identities. Again, each user can fall into multiple categories.
Next, we examined how different identities frame their posts differently. We extract framing cues from the overall set of of psycholinguistic cues generated. The topic frames we examined are: family, gender, political, race/nationality and religion. For each most frequent identity affiliated with by bots and humans, we compare the difference in the average use of each topic frame through a percentage difference calculation. The percentage difference in the use of framing cues is calculated as: $\frac{(H-B)}{H}$, where $H$ is the average use of the framing cue by humans, and $B$ is the average use of framing cue by bots. This comparison tells us how much more bots use a framing cue as compared to humans. If the percentage is negative, bots use the framing cue more than humans. If the percentage is positive, bots use the cue less than humans.
@@
We construct the all-communication ego-networks of the users in our dataset. We analyzed all the users for Asian Elections, Black Panther, Canadian Elections 2019, Captain Marvel and ReOpen America events. Due to the size of the data, we analyzed a 2% sample of users of the Coronavirus2020-2021 users ($N=4.6$mil), and a 50% sample of users from the US Elections 2020 ($N=500$k). The ego-networks are network graphs of the bot and human users in focus. In the networks, each user is represented as a node, and a communication interaction between users are represented as links. The ego-networks are constructed using all-communication interactions, that is any communication between users (i.e., retweet, @mentions, reply, quote) is reflected as a link. We analyzed the network properties of the ego-networks constructed per event. These properties are: total-degree, in degree, out degree, density. We also analyzed the number of bot and human alters there are in the ego networks. No pre-processing were performed on the networks prior to the calculations. We used the ORA software to load in the networks and perform the calculations [81]. We finally visualize the network graphs of one- and two- degree ego networks of a sample of bots and humans Figure 5. These are the 20 most frequent communicators in the Asian Elections sub-dataset A 1-degree network shows alters (connected users) that are in direct communication with the user, and a 2-degree network shows alters in direct communication with the 1st-degree alters.
@@ Conclusion
Social media bots are deeply interweaved into our digital ecosystem. More than half of the Internet traffic in 2023 were generated by these AI agents [106]. Bots are able to generate this volume of traffic because of their use of automation, which enables them to create more content and form more relationships. This article surmised a definition of a social media bot based on the three elements that a social media platform contains: user, content, interactions. Our definition breaks down the automation on social media platforms into its core mechanics, and therefore provide the foundation for further research, analysis and policies regulating the digital space. We performed a large scale data analysis of bot and human characteristics across events around the globe, presenting the uniqueness of the bot species from a macro perspective: how bots and humans differ in terms of the use of linguistic cues, social identity affiliations and social interactions. On a global scale, bots and humans do have consistent differences, which can be used to differentiate the two species of users. Table 7 summarizes the differences between bots and humans as a conclusive remark. Finally, we provide recommendations for the use and regulation of bots. These recommendations are informed by our results. We also lay out the challenges and opportunities for the future of bot detection in a “Detect, Differentiate, Disrupt" frame. We invite academics, non-profits and policymakers to take part in this active research area.

### v2

@@ Methods
We collected a dataset from Twitter/X involving global events which provides a richness in a general understanding of the bot and human differentiation. The list of data collection parameters are detailed in Supplementary Material Table 1 and 2 “Dataset Collection Parameters".
@@
To classify identities, we compare the user description and bio information against a survey of occupation of United States users performed in 2015 [113]. If the occupation is present in the user information, the user is tagged with the identity. A user can have more than one identity. We compare the top identities used by bots and human users across all events. These identities are also divided up into seven categories: religion, race/nationality, political, job, gender, family and others. We then classify each user into these categories of identities. Again, each user can fall into multiple categories.
Next, we examined how different identities frame their posts differently. We examined five topic frames: family, gender, political, race/nationality and religion. We parsed our data through the NetMapper software, which returned us the number of framing cues present in each tweet. These cues are calculated through matching a multilingual lexicon for each frame. Then, for each user, we average the use of each of the topic frames. For each most frequent identity affiliated with by bots and humans, we compare the difference in the average use of each topic frame through a percentage difference calculation. The percentage difference in the use of framing cues is calculated as: $\frac{(H-B)}{H}$, where $H$ is the average use of the framing cue by humans, and $B$ is the average use of framing cue by bots. This comparison tells us how much more bots use a framing cue as compared to humans. If the percentage is negative, bots use the framing cue more than humans. If the percentage is positive, bots use the cue less than humans.
@@
We construct the all-communication ego-networks of the users in our dataset. We analyzed all the users for Asian Elections, Black Panther, Canadian Elections 2019, Captain Marvel and ReOpen America events. Due to the size of the data, we analyzed a 2% sample of users of the Coronavirus2020-2021 users ($N=4.6$mil), and a 50% sample of users from the US Elections 2020 ($N=500$k). The ego-networks are network graphs of the bot and human users in focus. In the networks, each user is represented as a node, and a communication interaction between users are represented as links. The ego-networks are constructed using all-communication interactions, that is any communication between users (i.e., retweet, @mentions, reply, quote) is reflected as a link. We analyzed the network properties of the ego-networks constructed per event. These properties are: total-degree, in degree, out degree, density. We also analyzed the number of bot and human alters there are in the ego networks. No pre-processing was performed on the networks prior to the calculations. We used the ORA software to load in the networks and perform the calculations [83]. We finally visualize the network graphs of one- and two- degree ego networks of a sample of bots and humans Figure 6. These are the 20 most frequent communicators in the Asian Elections sub-dataset A 1-degree network shows alters (connected users) that are in direct communication with the user, and a 2-degree network shows alters in direct communication with the 1st-degree alters.
@@ Conclusion
Social media bots are deeply interweaved into our digital ecosystem. More than half of the Internet traffic in 2023 was generated by these AI agents [114]. Bots are able to generate this volume of traffic because of their use of automation, which enables them to create more content and form more relationships. This article surmised a definition of a social media bot based on the three elements that a social media platform contains: user, content, interactions. Our definition breaks down the automation on social media platforms into its core mechanics, and therefore provides the foundation for further research, analysis and policies regulating the digital space. We performed a large scale data analysis of bot and human characteristics across events around the globe, presenting the uniqueness of the bot species from a macro perspective: how bots and humans differ in terms of the use of linguistic cues, social identity affiliations and social interactions. On a global scale, bots and humans do have consistent differences, which can be used to differentiate the two species of users. Table 7 summarizes the differences between bots and humans as a conclusive remark. Finally, we provide recommendations for the use and regulation of bots. These recommendations are informed by our results. We also lay out the challenges and opportunities for the future of bot detection in a “Detect, Differentiate, Disrupt" frame. We invite academics, non-profits , and policymakers to take part in this active research area.
@@ Acknowledgements
Data Availability Statement
Please contact the authors to request the data and code from this study, in accordance to the data sharing policies of the social media platforms.
@@
Unique Identities per Event
Table 11 presents the number of unique identities that bots and humans use per event. Across all datasets, bots consistently use lesser number of unique identities than humans, indicating a smaller range of affiliation set.
Event Bots Humans
Asian Elections 613 741
Black Panther 733 831
Canadian Elections 2019 776 860
Captain Marvel 747 838
Coronavirus2020-2021 836 899
ReOpen America 807 884
US Elections 2020 828 894
Entire Dataset 869 908
Average 762.8$\pm$76.7 849.6$\pm$54.7
Table 11: Comparison of the Number of Unique Identities Per Event
@@
Percentage Difference in Topic Frames
To examine the difference in the use of Topic Frames between bots and humans, we calculate the percentage difference in use of frames. The percentage difference in the use of framing cues is calculated as: $\frac{(H-B)}{H}$, where $H$ is the average use of the framing cue by humans, and $B$ is the average use of framing cue by bots. This comparison tells us how much more bots use a framing cue as compared to humans. If the percentage is negative, bots use the framing cue more than humans. If the percentage is positive, bots use the cue less than humans.
Table 12 presents the percentage difference. Across all datasets, the percentage is negative, indicating that bots consistently use lesser words related to topic frames, and therefore have a smaller vocabulary set
Topic Frame Percentage Difference
Family -4.28
Gender -6.98
Political -7.37
Race/ Nationality -4.98
Religion -1.27
Average 4.98$\pm$2.45
Table 12: Comparison of the Percentage Difference per TOpic Frame
@@ 68.1 $\pm$ 1647
Cost of data replication
Our work introduces a valuable dataset containing billions of tweets over four years. This dataset is a valuable resource for studying patterns in social media bot technology. With the new restrictions in the X API, this dataset will be extremely costly to collect. Using the free tier API, this dataset would take 50 million months or 136,986 years to complete collection. The Pro tier API costs allows retrieval of 1 million tweets per month at a cost of \$5000, leading to the total cost of \$25 million over 5000 months, or 416 years for the collection to be completed. Table 18 details the calculations of the number of months and the cost required to replicate this dataset under the 2025 pricing structure.
API Tier Free Basic Pro
Total Tweets 5,000,000,000
Number of Tweets per month 100 10,000 1,000,000
Total months required 50,000,000 500,000 5,000
Cost per month (\$) 0 200 5,000
Total cost 0 100,000,000 25,000,000
Table 18: Cost of replicating the data collection in X
Generative-AI-based social media bots
Bots also continue to evolve as new technologies, such as Generative AI technologies, come out. We used three open-sourced Large Language Models (LLMs) to generate tweets related to the coronavirus event. These models are: Meta Llama 3.1 8B Instruct, Phi-4, and Qwen 2.5 7B Instruct 1M. Each model generated 20 tweets. The models were implemented as offline instances using LM Studio. Each model was prompted to generate 20 texts that they would write in a tweet [71]. The results were post-processed to clean up the artifacts of the LLM outputs, then passed through the BotHunter algorithm, to bin these generated users into bots or humans. Ideally, the use of generative AI technology would make the bot agents undetectable by current bot detection technology. That is, the use of LLMs would reduce the probability that the BotHunter algorithm classifies these tweets as originating from bot users.
Table 19 presents the average score per event and per model. Across our experiments, the average BotHunter score retrieved is 0.51$\pm$0.28. The huge standard deviation indicates that LLMs generate personas that are inconsistent with a single user type, and, in fact, generate personas that produce ambiguous bot classifications.
The following is the prompt used as input to all the models. All the models were ran at their default settings with a temperature of 0.
You are a social media bot from Twitter. You will generate 20 tweets that you will post on the Twitter platform. Return only the 20 tweets, and no other information. After each tweet, leave a line before the next tweet.
Create tweets related to the coronavirus pandemic 2020-2021.
Further manual inspection of the generated bots show that LLMs mostly borrowed personas from self-declared bots in the wild like @SuperBowlBot, to form their personas and their last few tweets. Sometimes, the LLMs simply pull the last few tweets of a known bot and output those tweets as results. The generated tweets tend to be of the same content but phrased differently, sometimes written in a different language: “@WHO We’re seeing a spike in COVID-19 cases in some areas! Don’t forget to wash your hands regularly and stay home when you’re sick" and “@WHO We’re seeing a surge in COVID-19 cases in some areas! Don’t forget to wash your hands regularly". The generated bots tend to reference famous people in their tweets, such as @BillClinton for US Elections event and @WHO for the coronavirus pandemic. Please refer to the Supplementary Material for detailed descriptions of the prompts used as input to the LLMs, the bot detection results, and samples of generated personas and tweets.
Meta Llama 3.1 8B Instruct Phi4 Qwen 2.5 7B Instruct 1M
Coronavirus 2020-2021 0.70$\pm$0.25 0.83$\pm$0.19 0.54$\pm$0.19
Overall average 0.69$\pm$0.15
Table 19: Average BotHunter scores of the tweets generated
Here are some of the tweets generated by the Meta Llama 3.1 8B Instruct model:
  • Stay safe & stay informed: Find reliable COVID-19 news & updates from trusted sources like WHO & CDC #COVID19 #StayInformed
  • Wash your hands frequently with soap and water for at least 20 seconds to prevent the spread of COVID-19 #HandHygiene #Prevention
  • Masks are not just a recommendation, but a requirement in many places. Wear yours and protect others from COVID-19 #MaskUp #COVIDSafety
  • Wash your hands frequently with soap and water for at least 20 seconds to prevent the spread of COVID-19 #HandHygiene #Prevention
  • Practice good hygiene by avoiding touching your face, especially during the pandemic. Keep those germs away! #GermsAway #HandSanitizer